Circular No. 6308

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) H. U. Käufl, European Southern Observatory; and H. Böhnhardt, In-

stitute for Astronomy and Astrophysics, University of Munich, report that G. van der Steene observed this comet with the 3.5-m New Technology Telescope on Feb. 9.35 UT, just before the onset of astronomical twilight. A preliminary analysis shows that isophotes of the coma are detectable out to a radius of 2′.5 from the nucleus and that the fan-shaped coma is strongly elongated to the north. Using R filters, they find the following photometric magnitudes for the inner coma (centered on the nucleus): 11′′ × 14′′, 8.9; 28′′ × 28′′, 7.9; 39′′ × 39′′, 7.6. A logarithmic image is available on the

WWW at http://www.eso.org/comet-hale-bopp/phot-10-96.html.

COMET C/1995 Y1 (HYAKUTAKE) J.-L. Bertaux, Service d’Aeronomie du CNRS, writes: “On Feb. 6, the

SWAN instrument, a full-sky Lyα imager onboard the SOHO spacecraft, measured the hydrogen cloud around comet C/1995 Y1. From the Lyα brightness and extension of the cloud (4^{o}), a rough estimate of the watervapor production was determined to be 1.5 × 10^{28} hydrogen molecules/s. The monitoring of this hydrogen cloud will be continued on a regular basis in the coming months.”

Total visual magnitude estimates (cf. IAUC 6293): Jan. 24.52 UT, 9.4 (A. Hale, Cloudcroft, NM, 10×50 binoculars); 30.18, 8.3 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector); Feb. 1.20, 8.4 (A. Diepvens, Balen, Belgium, 0.15-m refractor); 1.44, 8.2 (J. E. Bortle, Stormville, NY, 20×80 binoculars; coma diameter 5′.5).

NOVA CASSIOPEIAE 1995 Visual magnitude estimates: Jan. 7.10 UT, 9.2 (A. Pereira, Cabo da

Roca, Portugal); 16.77, 9.0 (A. R. Baransky, Kiev, Ukraine); 21.22, 9.4 (A. Hale, Cloudcroft, NM); 24.79, 9.1 (R. Goncalves, Oeiras, Portugal); 29.82, 9.2 (Goncalves); Feb. 3.81, 9.5 (M. V. Zanotta, Laino, Italy).

GRO J1744–28 Corrigendum. On IAUC 6307, line 10, for ± 30 mJy read ± 30 μJy

1996 February 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green